OpenStudy (anonymous):

find the area of a triangle with a base 2 1/5yards and height 5/8 yard. (Hint the area of a triangle equals half the product of its base and height.)

OpenStudy (zale101):

Area= 1/2 base x height

OpenStudy (zale101):

Plug in the base and the height, and Solve!

OpenStudy (anonymous):

@Zale101 10/1+5/8 IM LOST

OpenStudy (zale101):

Or... \[Area= (2 \frac{ 1 }{ 5 }* \frac{ 5 }{ 8 })/(2)\]

OpenStudy (anonymous):

IM STILL STUCK @Zale101 IS IT 11/5*5/8 =11/5*8/5/2=7.04

OpenStudy (zale101):

No, it's Area= (11/5) * (5/8)/2 = (55/40)*(1/2)=55/80= 11/16

OpenStudy (anonymous):

WOW i didnt even think of doing it tht way but thanks 4 ya help @Zale101

OpenStudy (zale101):

np